Find the mean by direct and short-cut method from the data showing marks of students in a class in an economics test; 40, 50, 55, 78, 58.


1
Expert's answer
2021-06-29T10:25:25-0400
"\frac{40+50+55+78+58}{5}=\frac{281}{5}=56.2"